# Eyes Can Deceive: Benchmarking Counterfactual Reasoning Capabilities of Multimodal Large Language Models

### v1

@@ Abstract.
Counterfactual reasoning, as a crucial manifestation of human intelligence, refers to making presuppositions based on established facts and extrapolating potential outcomes. Existing multimodal large language models (MLLMs) have exhibited impressive cognitive and reasoning capabilities, which have been examined across a wide range of Visual Question Answering (VQA) benchmarks. Nevertheless, how will existing MLLMs perform when faced with counterfactual questions? To answer this question, we first curate a novel CounterFactual MultiModal reasoning benchmark, abbreviated as CFMM, to systematically assess the counterfactual reasoning capabilities of MLLMs. Our CFMM comprises six challenging tasks, each including hundreds of carefully human-labeled counterfactual questions, to evaluate MLLM’s counterfactual reasoning capabilities across diverse aspects. Through experiments, interestingly, we find that existing MLLMs prefer to believe what they see, but ignore the counterfactual presuppositions presented in the question, thereby leading to inaccurate responses. Furthermore, we evaluate a wide range of prevalent MLLMs on our proposed CFMM. The significant gap between their performance on our CFMM and that on several VQA benchmarks indicates that there is still considerable room for improvement in existing MLLMs toward approaching human-level intelligence. On the other hand, through boosting MLLMs performances on our CFMM in the future, potential avenues toward developing MLLMs with advanced intelligence can be explored.
Counterfactual reasoning, Multimodal large language model, Visual question answering, New benchmark
ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
1. Introduction
@@
Figure 1. The comparison of our proposed CFMM with existing VQA benchmarks in terms of dataset scale and the degree of approaching the human-level intelligence. We categorize all VQA benchmarks into three groups based on their proximity to human-level intelligence and distinguish between these groups using different markers. We also provide examples within each group for intuitive understanding.
@@
Recently, the Multimodal Large Language Model (MLLM) (Achiam et al., 2023; Alayrac et al., 2022; Bai et al., 2023; Dai et al., 2024; Li et al., 2023b, c; Peng et al., 2023; Jiao et al., 2024; Zhu et al., 2023) has been a rising research hotspot due to its remarkable content comprehension and logic reasoning capabilities. Harnessing the Large Language Model (LLM) as the brain, MLLM bridges the visual world and language space, empowering the model to observe the vast world and enrich its mastered knowledge. Pioneering works like Qwen-VL (Bai et al., 2023), LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) pursue GPT-4V-like (Achiam et al., 2023) visual perception and comprehension abilities by performing multimodal instruction tuning. Subsequent approaches, including Kosmos-2 (Peng et al., 2023), Shikra (Chen et al., 2023), Lumen (Jiao et al., 2024), etc., further enhance the visual comprehension by endowing the model with fine-grained grounding capabilities.
@@
As diverse MLLMs continue to rapidly evolve, benchmarks evaluating the versatile emergent capabilities of MLLMs are concurrently being developed to keep pace with their advancements. Early benchmarks like VQAv2 (Goyal et al., 2017), COCO-Caption (Lin et al., 2014), GQA (Hudson and Manning, 2019), OK-VQA (Marino et al., 2019), etc., focus on assessing visual content comprehension capabilities on specific domains. In addition, the evaluation metrics they employed can not comprehensively reflect model performances. To sufficiently assess the emergent capabilities of existing MLLMs, a new generation of benchmarks (Fu et al., 2023a; Li et al., 2023a; Liu et al., 2023) has been introduced. These benchmarks encompass a broader range of questions and employ more reliable evaluation metrics. A key advancement of these benchmarks is their emphasis on requiring the synergy of vision and language modalities for more complex reasoning to derive answers. However, the reasoning capabilities they addressed still lag far behind real human intelligence.
Figure 2. Performances of prevalent MLLMs achieved on existing VQA benchmarks and our proposed CFMM benchmark. The evident performance gap between prevalent VQA benchmarks and our proposed CFMM indicates that existing MLLMs struggle to address counterfactual reasoning.
@@
One of the most pivotal aspects of human intelligence lies in its capacity for imagination. Building upon the perceived world and mastered knowledge, imagination empowers humans to explore hypothetical scenarios and potential outcomes beyond the confines of reality. This cognitive process is referred to as the “counterfactual reasoning” in psychology (Epstude and Roese, 2008; Van Hoeck et al., 2015). Essentially, a counterfactual question can be formulated as: “Given F, if P happens, how will Q unfold?”. Here, F represents the known facts, P indicates the imaginary presupposition of changing the status of F, and Q refers to queried concepts like numbers, status, actions, etc. In the scenario of multimodal content comprehension, images are generally employed as F to reflect the real world, while P and Q are expressed with natural language. Interestingly, we observe that Although MLLMs demonstrate strong performance on prevalent VQA benchmarks, they still struggle to tackle these counterfactual questions as shown in Fig 2. Through further analysis, we find that existing MLLMs are prone to believe what they see, but are lazy to comprehend the presupposition presented in sentences, thus leading to incorrect responses. These results indicate that there is still significant room for improvement for MLLMs to approach the level of human intelligence.
@@
To assess the advanced counterfactual reasoning capabilities of existing MLLMs and explore potential avenues for further optimizing MLLMs, we introduce CFMM, a CounterFactual MultiModal reasoning benchmark in this paper. Our CFMM comprises a suite of 6 tasks to comprehensively evaluate model’s counterfactual reasoning capabilities in diverse aspects. As shown in Fig 1, compared with prevalent VQA benchmarks, our CFMM presents higher demands on the MLLM’s intelligence and encompasses substantial amounts of data samples. As for the dataset curation, firstly, to avoid the data leakage problem, we collect 1,200 images from the validation set of COCO-2014 (Lin et al., 2014) as visual resources. For each image, we resort to crowd-source to manually label paired questions with one requiring counterfactual reasoning and another that does not, aiming to investigate whether the model genuinely comprehends the counterfactual presupposition. In order to obtain objective and accurate evaluation metrics, the answers are presented as binary choices, and accuracy serves as the evaluation metric. To eliminate the model’s bias toward preferring a particular choice position, we follow the methodology from previous work (Li et al., 2023a) by adopting the answer ranking strategy for evaluation. Specifically, conditioned on embedded image tokens and the query sentence semantics, we select the choice with highest probability predicted by the MLLM. In addition, we extensively evaluate the prevalent MLLMs on our CFMM dataset and provide further in-depth analyses.
@@
  • We curate a challenging counterfactual dataset, CFMM, which is capable of comprehensively evaluating current MLLMs on counterfactual tasks in six categories: count, color, size, shape, direction, and common sense.
@@
  • Using the answer ranking strategy for evaluation, we provide an objective and accurate evaluation for current state-of-the-art MLLMs, experiment results indicate significant performance degradation for all MLLMs when confronted with counterfactual tasks.
@@
  • We analyze the performance of current state-of-the-art MLLMs on the CFMM dataset and specifically discuss how lightweight techniques such as few-shot learning, ICL, and CoT enhance the model’s counterfactual reasoning capabilities, in the hope of providing meaningful support for further research in these area.
2. Related Works
2.1. Multimodal Large Language Models
@@
With the scaling of data size and model size, large language models (LLMs) have shown remarkable performance across various linguistic tasks (Zhang et al., 2022a; Chung et al., 2022; Touvron et al., 2023; Anil et al., 2023). Researchers have delved into extensive explorations to extend the robust generalization capabilities inherent in LLMs into multimodal counterparts (MLLMs). Flamingo (Alayrac et al., 2022) and Open-Flamingo (Awadalla et al., 2023) incorporate two trainable modules, namely the Perceiver Resampler and XAttn-Dense, which enhance the few-shot performance of the model. LLaMA-Adapter (Zhang et al., 2023a; Gao et al., 2023) and Otter (Li et al., 2023f) insert trainable cross-attention layers into LLMs to integrate visual and language information. BLIP2 (Li et al., 2023b) proposes Q-Former to bridge the modality gap between the frozen LLM and vision encoder. InstructBLIP (Dai et al., 2024), MiniGPT-4 (Zhu et al., 2023), VPGTrans (Zhang et al., 2024), mPLUG-OWL (Ye et al., 2023) retain Q-Former while utilizing manually-annotated or GPT4-generated high-quality data for second-stage fine-tuning to enhance the models’ instruction-following abilities. LLaVA (Liu et al., 2024) directly maps visual features to LLM through a lightweight projection layer. Shikra (Chen et al., 2023) and Kosmos-2 (Peng et al., 2023) further enhance MLLMs’ visual grounding abilities such as region description and localization. Besides, Qwen-VL (Bai et al., 2023) increase the input resolution of the visual encoder to 448 × 448, reducing the information loss caused by down-sampling. Monkey (Li et al., 2023c) goes even further by splitting high-resolution images into smaller patches, enabling the model to support resolutions up to 1344×896. In addition, the closed-source business model GPT-4v (Achiam et al., 2023) also demonstrates remarkable image comprehension capabilities. In CFMM-Bench, we provide a comprehensive evaluation of these models to thoroughly assess their multimodal counterfactual reasoning capabilities.
2.2. Counterfactual Reasoning
@@
”Counterfactual” is an informal term that refers to different perturbations made to factual information. There have been several attempts to assess LLMs’ counterfactual reasoning abilities. CRASS (Frohberg and Binder, 2022) proposes the first question answering dataset that requires LLMs to understand the ”if” clause consisting of a counterfactual presupposition. IfQA (Yu et al., 2023) explores hypothetical changes of world knowledge in open-domain question answering scenarios. (Li et al., 2023d) further investigates the impact of cue words and whether the model could infer real-world states from counterfactual conditions. (Wu et al., 2023b) designs counterfactual counterparts for 11 default tasks, and evaluating recent strong LLMs on them. ACQUIRED (Wu et al., 2023a) proposes a counterfactual version of video question answering, where the model is required to infer events that did not actually occur in the video based on the hypothetical language descriptions. VQA-CP (Agrawal et al., 2018) and CF-VQA (Niu et al., 2021) define counterfactual visual question answering as situations where the attribute distributions differ between the training and testing sets. Most related to our work is C-VQA (Zhang et al., 2023b), which leverages Chat-GPT to incorporate counterfactual presuppositions into questions sourced from VQAv2 (Goyal et al., 2017). However, compared to CFMM, the ”if” conditions in C-VQA only involve changes in quantity, lacking diversity in question types.
2.3. In-Context Learning
@@
LLMs exhibit promising in-context learning (ICL) capabilities (Brown et al., 2020), that is, learning from a few input-label examples from downstream tasks without training. In terms of in-context examples selection, KATE (Liu et al., 2021) retrieves the k-nearest neighbors that are semantically-similar to a test sample. EPR (Min et al., 2022) extends this approach by training a prompt retriever with contrastive loss, which then select demonstrations based on the confidence scores during inference. Auto-CoT (Zhang et al., 2022b) categorizes problems into clusters, employing zero-shot CoT to automatically construct demonstrations for each cluster. Multimodal ICL (M-ICL) is a complex and relatively unexplored field. PICa (Yang et al., 2022) directly translates images into textual descriptions as inputs to GPT-3. MM-React (Yang et al., 2023) additionally incorporates outputs from various visual experts (e.g. dense captioning, object tagging) as inputs to Chat-GPT. Science-QA (Lu et al., 2022), MMIC-IT (Li et al., 2023e) and MMICL (Zhao et al., 2023) directly provide demonstrations within the dataset, allowing the model to utilize instruction-tuning to further enhance the capabilities of M-ICL. In this paper, we validated the effectiveness of the in-context learning ability of MLLMs on CFMM to a certain extent.
3. CFMM Benchmark
@@
Figure 6. Demonstration of questions of six types included in our CFMM. For each image, we provide the corresponding basic and counterfactual questions in the first two rows, as well as the answer options in the last row. We include the correct answer of each question in the bracket at the beginning of the question sentence.Note that counterfactual presuppositions may also serve as distractors irrelevant to answer selection as shown in the left case in the last row.
@@
We present the CounterFactual MultiModal reasoning (CFMM) benchmark, a manually curated benchmark to evaluate the counterfactual reasoning capabilities of MLLMs. In this section, we will first elaborate on the definition and taxonomy of counterfactual questions, and then introduce the dataset construction process, as well as evaluation metrics calculation. Additionally, we thoroughly explore and analyze the statistical insights within the CFMM-bench.
3.1. Problem Definition
@@
Counterfactual questions refer to new questions formed by adding imaginary presuppositions to known facts of basic questions. In the NLP field, ”facts” are often declarative statements describing background information, ”presuppositions” are hypothetical alterations to this information, and ”questions” often involve content related to causal inferences. This paradigm extends into the multimodal domain, where ”facts” are substituted with multimodal elements like images or videos. For MLLM, answering counterfactual questions essentially requires conditioning on presuppositions and facts to output the answer with the highest probability. We formulate it as a function $f$: $X$$Y$ that maps the input $x\in X$ to the output $y\in Y$:
$f(v,w_{cf},w_{q})=\mathop{\arg\max}\limits_{y^{\prime}}\mathbbm{P}(y^{\prime}|v,w_{cf},w_{q}).$
@@
Here, $y^{\prime}$ is the output of MLLM obtained through an appropriate decoder. $v$, and $w_{cf}$, $w_{q}$ represent the image, imaginary presuppositions, and basic question respectively.
3.2. Diverse Aspects of Counterfactual Abilities
  • Common: Change world conditions and asking the model to reason based on external knowledge. This can encompass a wide range of questions that require the model to apply logic and knowledge to new scenarios.
@@ 3.2. Diverse Aspects of Counterfactual Abilities
C-VQA VQAv2 Automatic 2 3144
CFMM COCO Human 6 2400
@@
Table 1. Statistical comparison with existing counterfactual reasoning benchmarks in both Text-QA and VQA tasks.
3.3. Dataset Construction
Data Source. The counterfactual question involves hypothetical conditions that entail changes in different aspects, which requires the images themselves to potentially incorporate such information. COCO(Lin et al., 2014) is a dataset tailored for object detection and image captioning. We choose to annotate images from it’s validation set due to the diverse object categories and rich visual scenes within.

### v3

@@ Abstract
Counterfactual reasoning, as a crucial manifestation of human intelligence, refers to making presuppositions based on established facts and extrapolating potential outcomes. Existing multimodal large language models (MLLMs) have exhibited impressive cognitive and reasoning capabilities, which have been examined across a wide range of Visual Question Answering (VQA) benchmarks. Nevertheless, how will existing MLLMs perform when faced with counterfactual questions? To answer this question, we first curate a novel CounterFactual MultiModal reasoning benchmark, abbreviated as CFMM, to systematically assess the counterfactual reasoning capabilities of MLLMs. Our CFMM comprises six challenging tasks, each including hundreds of carefully human-labeled and GPT-generated counterfactual questions, to evaluate MLLM’s counterfactual reasoning capabilities across diverse aspects. Through experiments, interestingly, we find that existing MLLMs prefer to believe what they see, but ignore the counterfactual presuppositions presented in the question, thereby leading to inaccurate responses. Furthermore, we evaluate a wide range of prevalent MLLMs on our proposed CFMM. The significant gap between their performance on our CFMM and that on several VQA benchmarks indicates that there is still considerable room for improvement in existing MLLMs toward approaching human-level intelligence. On the other hand, through boosting MLLMs performances on our CFMM in the future, potential avenues toward developing MLLMs with advanced intelligence can be explored.
1 Introduction
@@
Figure 1: The comparison of our proposed CFMM with existing VQA benchmarks in terms of dataset scale and the degree of approaching the human-level intelligence. We categorize all VQA benchmarks into three groups based on their proximity to human-level intelligence and distinguish between these groups using different markers. We also provide examples within each group for intuitive understanding.
@@
Recently, the Multimodal Large Language Models (MLLMs) (Achiam et al. 2023; Alayrac et al. 2022; Bai et al. 2023; Dai et al. 2024; Li et al. 2023d, e; Peng et al. 2023; Jiao et al. 2024; Zhu et al. 2023) have been a rising research hotspot due to their remarkable content comprehension and logic reasoning capabilities. Harnessing the Large Language Model (LLM) as the brain, MLLM bridges the visual world and language space, empowering the model to observe the vast world and enrich its mastered knowledge. Pioneering works like Qwen-VL (Bai et al. 2023), LLaVA (Liu et al. 2024) and MiniGPT-4 (Zhu et al. 2023) pursue GPT-4V-like (Achiam et al. 2023) visual perception and comprehension abilities by performing multimodal instruction tuning. Subsequent approaches, including Kosmos-2 (Peng et al. 2023), Shikra (Chen et al. 2023), Lumen (Jiao et al. 2024), etc., further enhance the visual comprehension by endowing the model with fine-grained grounding capabilities.
@@
As diverse MLLMs continue to rapidly evolve, benchmarks evaluating the versatile emergent capabilities of MLLMs are concurrently being developed to keep pace with their advancements. Early benchmarks like VQAv2 (Goyal et al. 2017), COCO-Caption (Lin et al. 2014), GQA (Hudson and Manning 2019), OK-VQA (Marino et al. 2019), etc., focus on assessing visual content comprehension capabilities on specific domains. In addition, the evaluation metrics they employed can not comprehensively reflect model performances. To sufficiently assess the emergent capabilities of existing MLLMs, a new generation of benchmarks (Fu et al. 2023a; Li et al. 2023a; Liu et al. 2023) has been introduced. These benchmarks encompass a broader range of questions and employ more reliable evaluation metrics. A key advancement of these benchmarks is their emphasis on requiring the synergy of vision and language modalities for more complex reasoning to derive answers. However, the reasoning capabilities they addressed still lag far behind real human intelligence.
@@
One of the most pivotal aspects of human intelligence lies in its capacity for imagination. Building upon the perceived world and mastered knowledge, imagination empowers humans to explore hypothetical scenarios and potential outcomes beyond the confines of reality. This cognitive process is referred to as the “counterfactual reasoning” in psychology (Epstude and Roese 2008; Van Hoeck, Watson, and Barbey 2015). Essentially, a counterfactual question can be formulated as: “Given F, if P happens, how will Q unfold?”. Here, F represents the known facts, P indicates the imaginary presupposition of changing the status of F, and Q refers to queried concepts like numbers, status, actions, etc. In the scenario of multimodal content comprehension, images are generally employed as F to reflect the real world, while P and Q are expressed with natural language. Interestingly, we observe that Although MLLMs demonstrate strong performance on prevalent VQA benchmarks, they still struggle to tackle these counterfactual questions. Through further analysis, we find that existing MLLMs are prone to believe what they see, but are lazy to comprehend the presupposition presented in sentences, thus leading to incorrect responses. These results indicate that there is still significant room for improvement for MLLMs to approach the level of human intelligence.
@@
To assess the advanced counterfactual reasoning capabilities of existing MLLMs and explore potential avenues for further optimizing MLLMs, we introduce CFMM, a CounterFactual MultiModal reasoning benchmark in this paper. Our CFMM comprises a suite of 6 tasks to comprehensively evaluate model’s counterfactual reasoning capabilities in diverse aspects. As shown in Fig 1, compared with prevalent VQA benchmarks, our CFMM presents higher demands on the MLLM’s intelligence and encompasses substantial amounts of data samples. As for the dataset curation, firstly, to avoid the data leakage problem, we collect 1,200 images from the validation set of COCO-2014 (Lin et al. 2014) as visual resources. For each image, we rely on crowd-sourcing to manually label paired questions, with one involving a counterfactual presupposition and the other without. Afterward, we leverage GPT-4V to generate three additional counterfactual questions for diversity. Statistically, we have a total of 6,000 questions, with one basic question and four counterfactual questions per image. In order to obtain objective and accurate evaluation metrics, the answers are presented as binary choices, and accuracy serves as the evaluation metric. To eliminate the model’s bias toward preferring a particular choice position, we follow the methodology from previous work (Li et al. 2023a) by adopting the answer ranking strategy for evaluation. Specifically, conditioned on embedded image tokens and the query sentence semantics, we select the choice with highest probability predicted by the MLLM. In addition, we extensively evaluate the prevalent MLLMs on our CFMM dataset and provide further in-depth analyses.
@@
  • We curate a novel counterfactual reasoning evaluation benchmark, CFMM, designed to comprehensively assess current MLLMs on counterfactual tasks across diverse aspects, including counting, color, size, shape, direction, and common sense.
@@
  • We provide an objective and accurate evaluation for current state-of-the-art MLLMs, experiment results indicate significant performance degradation for all MLLMs when confronted with counterfactual tasks.
@@
  • We analyze the performance of current state-of-the-art MLLMs on the CFMM dataset and specifically discuss how lightweight techniques such as few-shot learning, ICL, and CoT affect the model’s counterfactual reasoning capabilities, in the hope of providing meaningful support for further research in these areas.
2 Related Works
@@
Multimodal Large Language Models. With the scaling of data size and model size, large language models (LLMs) have shown remarkable performance across various linguistic tasks (Zhang et al. 2022a; Chung et al. 2022; Touvron et al. 2023; Anil et al. 2023). Researchers have delved into extensive explorations to extend the robust generalization capabilities inherent in LLMs into multimodal counterparts (MLLMs). Flamingo (Alayrac et al. 2022) and Open-Flamingo (Awadalla et al. 2023) incorporate two trainable modules, namely the Perceiver Resampler and XAttn-Dense, which enhance the few-shot performance of the model. LLaMA-Adapter (Zhang et al. 2023b; Gao et al. 2023) and Otter (Li et al. 2023c) insert trainable cross-attention layers into LLMs to integrate visual and language information. BLIP2 (Li et al. 2023d) proposes Q-Former to bridge the modality gap between the frozen LLM and vision encoder. InstructBLIP (Dai et al. 2024), MiniGPT-4 (Zhu et al. 2023), VPGTrans (Zhang et al. 2024), mPLUG-OWL (Ye et al. 2023) retain Q-Former while utilizing manually-annotated or GPT4-generated high-quality data for second-stage fine-tuning to enhance the models’ instruction-following abilities. LLaVA (Liu et al. 2024) directly maps visual features to LLM through a lightweight projection layer. Shikra (Chen et al. 2023) and Kosmos-2 (Peng et al. 2023) further enhance MLLMs’ visual grounding abilities such as region description and localization. Besides, Qwen-VL (Bai et al. 2023) increase the input resolution of the visual encoder to 448 × 448, reducing the information loss caused by down-sampling. Monkey (Li et al. 2023e) goes even further by splitting high-resolution images into smaller patches, enabling the model to support resolutions up to 1344×896. In addition, the closed-source business model GPT-4V (Achiam et al. 2023) also demonstrates remarkable image comprehension capabilities. In CFMM, we provide a comprehensive evaluation of these models to thoroughly assess their multimodal counterfactual reasoning capabilities.
@@
Counterfactual Reasoning. “Counterfactual” is an informal term that refers to different perturbations made to factual information. There have been several attempts to assess LLMs’ counterfactual reasoning abilities. CRASS (Frohberg and Binder 2022) proposes the first question answering dataset that requires LLMs to understand the “if” clause consisting of a counterfactual presupposition. IfQA (Yu et al. 2023) explores hypothetical changes of world knowledge in open-domain question answering scenarios. (Li, Yu, and Ettinger 2023) further investigates the impact of cue words and whether the model could infer real-world states from counterfactual conditions. (Wu et al. 2023b) designs counterfactual counterparts for 11 default tasks, and evaluating recent strong LLMs on them. ACQUIRED (Wu et al. 2023a) proposes a counterfactual version of video question answering, where the model is required to infer events that did not actually occur in the video based on the hypothetical language descriptions. VQA-CP (Agrawal et al. 2018) and CF-VQA (Niu et al. 2021) define counterfactual visual question answering as situations where the attribute distributions differ between the training and testing sets. Most related to our work is C-VQA (Zhang et al. 2023a), which leverages ChatGPT to incorporate counterfactual presuppositions into questions sourced from VQAv2 (Goyal et al. 2017). However, constrained by the questions of VQA-v2 and their automated annotation process, its questions conditions lack diversity in question types. Additionally, the answer types are limited to numerical and boolean, which may not fully capture performance in open-world scenarios.
@@
In-Context Learning. LLMs exhibit promising in-context learning (ICL) capabilities (Brown et al. 2020), that is, learning from a few input-label examples from downstream tasks without training. In terms of in-context examples selection, KATE (Liu et al. 2021) retrieves the k-nearest neighbors that are semantically-similar to a test sample. EPR (Min et al. 2022) extends this approach by training a prompt retriever with contrastive loss, which then select demonstrations based on the confidence scores during inference. Auto-CoT (Zhang et al. 2022b) categorizes problems into clusters, employing zero-shot CoT to automatically construct demonstrations for each cluster. Multimodal ICL (M-ICL) is a complex and relatively unexplored field. PICa (Yang et al. 2022) directly translates images into textual descriptions as inputs to GPT-3. MM-React (Yang et al. 2023) additionally incorporates outputs from various visual experts (e.g., dense captioning, object tagging) as inputs to ChatGPT. Science-QA (Lu et al. 2022), MMIC-IT (Li et al. 2023b) and MMICL (Zhao et al. 2023) directly provide demonstrations within the dataset, allowing the model to utilize instruction-tuning to further enhance the capabilities of M-ICL. In this paper, we validated the effectiveness of the in-context learning ability of MLLMs on CFMM to a certain extent.
3 CFMM Benchmark
@@
Figure 2: Demonstration of questions of six types included in our CFMM. For each image, we provide the corresponding basic and counterfactual questions in the first two rows, as well as the answer options in the last row. We include the correct answer of each question in the bracket at the beginning of the question sentence. Note that counterfactual presuppositions may also serve as distractors irrelevant to answer selection as shown in the left case in the last row.
@@
We present the CounterFactual MultiModal reasoning (CFMM) benchmark, a manually curated benchmark to evaluate the counterfactual reasoning capabilities of MLLMs. In this section, we will first elaborate on the definition and taxonomy of counterfactual questions, and then introduce the dataset construction process, as well as evaluation metrics calculation.
3.1 Problem Definition
@@
Counterfactual questions refer to new questions formed by adding imaginary presuppositions to known facts of basic questions. In the NLP field, “facts” are often declarative statements describing background information, “presuppositions” are hypothetical alterations to this information, and “questions” often involve content related to causal inferences. This paradigm extends into the multimodal domain, where “facts” are substituted with multimodal elements like images or videos. For MLLM, answering counterfactual questions essentially requires conditioning on presuppositions and facts to output the answer with the highest probability. We formulate it as a function $f$: $X$$Y$ that maps the input $x\in X$ to the output $y\in Y$:
$f(v,w_{cf},w_{q})=\mathop{\arg\max}\limits_{y^{\prime}}\mathbf{P}(y^{\prime}|v,w_{cf},w_{q}).$
@@
Here, $y^{\prime}$ is the output of MLLM obtained through an appropriate decoder. $v$ and $w_{cf}$, $w_{q}$ represent the image, imaginary presuppositions, and basic question respectively.
@@
C-VQA-Real VQAv2 Automatic 2 3144
CFMM COCO Human & Auto. 6 6000
@@
Table 1: Statistical comparison with existing counterfactual reasoning benchmarks in both Text-QA and VQA tasks.
3.2 Diverse Aspects of Counterfactual Abilities
  • Common: Change world conditions and ask the model to reason based on external knowledge. This can encompass a wide range of questions that require the model to apply logic and knowledge to new scenarios.
3.3 Dataset Construction
Data Source. Counterfactual question involves hypothetical conditions that entail changes in different aspects, which requires images to contain rich scenarios and potential semantic information for hypothesis. The COCO dataset (Lin et al. 2014) is widely recognized for its extensive image quantity and its focus on object detection, image captioning, and scene understanding. We chose to annotate images from its validation set because of the diverse object categories and rich visual scenes it offers, which capture complex everyday scenarios closely aligned with real-world distributions. This makes COCO particularly well-suited for assessing models’ multimodal counterfactual reasoning capabilities in open-world scenarios. While our images are sourced from COCO, all questions and answers are all newly annotated, ensuring the fairness of our benchmark as the models’ training won’t include these question-answer pairs, even if they had seen the COCO images before.